# Thermoelectric cooling of a finite reservoir coupled to a quantum dot

### v1

@@ Abstract
We investigate non-equilibrium transport of charge and heat through an interacting quantum dot coupled to a finite electron reservoir. Both the quantum dot and the finite reservoir are coupled to conventional electric contacts, i.e., infinite electron reservoirs, between which a bias voltage can be applied. We develop a phenomenological description of the system, combining a rate equation for transport through the quantum dot with standard linear response expressions for transport between the finite and infinite reservoirs. The finite reservoir is assumed to be in a quasi-equilibrium state with time-dependent chemical potential and temperature which we solve for self-consistently. We show that the finite reservoir can have a large impact on the stationary state transport properties, including a shift and broadening of the Coulomb diamond edges. We also demonstrate that there is a region around the conductance lines where a heat current flows out of the finite reservoir. Our results reveal the dependence of the temperature that can be reached by this thermoelectric cooling on the system parameters, in particular the coupling between the finite and infinite reservoirs and additional heat currents induced by electron-phonon couplings, and can thus serve as a guide to experiments on quantum dot-enabled thermoelectric cooling of finite electron reservoirs. Finally, we study the full dynamics of the system, with a particular focus on the timescales involved in the thermoelectric cooling.
@@ I Introduction
Experimental advances in the fabrication of nanoscale electronic devices have made it possible to employ measurements of non-equilibrium transport to deduce the properties of a (quantum) system. Such devices include a wide range of systems, e.g., molecular junctions, carbon nanotubes, graphene structures and quantum dots (QDs)[1, 2, 3, 4, 5]. They are a fundamental building block for applications such as transistors, (thermal) diodes, nanoscale heat engines, and sensors [6, 7, 8, 9, 10, 11, 12]. From a theoretical point of view, the description of open quantum systems out of equilibrium remains a challenging task [13]. In non-equilibrium transport scenarios, it is common to assume that the system is in contact with infinite reservoirs at different temperatures and chemical potentials. Although this is a suitable description for a number of experimental situations, accounting for the finite aspect of reservoirs is needed in many nanoscale, non-equilibrium scenarios [14]. Various experiments on, e.g., superconducting [15, 16] or ultracold atom platforms [17] demonstrate intricate control over the non-equilibrium properties of the system, including access to the state of the reservoir [18, 19].
@@
Particle transport also induces the exchange of energy and heat between different parts of a system. It is natural to consider thermal properties and thermoelectric behavior of finite reservoirs, which are the subject of ongoing interest [33]. Generally, thermoelectric effects describe a system’s or material’s ability to convert a temperature difference to an electric voltage (or vice-versa). A high thermoelectric efficiency results from strong energy asymmetries in electron transport properties  [34, 35], which result from sharp resonances in meso- and nanoscale systems  [36, 37]. They readily occur in QD devices, promoting QDs to an ideal platform for experimental studies of thermoelectric behavior in quantum devices, see e.g. [38, 39, 40, 41, 12, 42]. In particular, an immediate use of thermoelectric properties is Peltier cooling, where a bias voltage can be used to force heat to flow from a cold to a hot reservoir. Such electronic cooling has been realized experimentally using QD systems [43], normal metal-insulator-superconductor junctions [44, 45] and has been investigated in cold atomic gases [46, 47]. In this context, given that any physical device or system being refrigerated is necessarily of a finite size, it is essential to understand transport and thermoelectric behaviors in non-equilibrium settings with finite reservoirs.
@@
The paper is organized as follows. In Section II, the system is introduced, and its non-equilibrium dynamics is modeled using a combination of a rate equation for the QD and a mesoscopic, linear-response approach for the interaction between the finite and infinite reservoirs. In Section III, we solve for the non-equilibrium stationary state and analyse the impact of the finite reservoir on the transport through the QD. The thermoelectric refrigeration of the finite reservoir is discussed in Section IV. Finally, we investigate the system’s full dynamics in Section V.
@@ II Model
The system consists of a single level, spin-degenerate QD tunnel-coupled to an infinite reservoir on the right and a finite reservoir on the left, see Fig. 1. The finite reservoir is additionally in contact with a second infinite reservoir to the left. All reservoirs are described by non-interacting electrons. For the finite reservoir we assume that the internal relaxation is fast compared to any other time-scale in the system, such that it is in a quasi-equilibrium state with a well-defined (but time-dependent) chemical potential $\mu_{L}(t)$ and temperature $T_{L}(t)$, and is described by the Fermi-Dirac distribution $f(E,\mu_{L}(t),T_{L}(t)$ with $f(E,\mu,T)~{}=~{}\{\exp\left[(E-\mu)/T\right]+1\}^{-1}$. Throughout the paper we set Boltzmann’s constant $k_{B}=1$, the elementary charge $e=1$ and $\hbar=1$.
@@
where $\Gamma_{L}$ is the tunneling rate between the QD and finite reservoir. The total rate $\Gamma_{L}^{\text{in}}\sim 2\Gamma_{L}$ in Eq. 2 takes into account the spin degeneracy of the state with one electron on the QD. For the exchange with the right reservoir the tunnel rates are time-independent and given by
@@
$J^{\text{QD}}_{L/R}(t)$ $=\varepsilon I^{\text{QD}}_{L/R}(t).$
@@
where $\tilde{T}(t)=(T_{L}(t)+\bar{T}_{L})/2$. In the strict linear response regime, Eq. 6 reduces to the Wiedemann-Franz law [50].
@@
Here we introduced the heat currents flowing into the finite reservoir
@@
$\dot{Q}_{L}(t)$ $=J_{L}(t)-\mu_{L}(t)I_{L}(t).$
This shows that the dynamics of the reservoir temperature $T_{L}(t)$ is governed by the heat flows into the reservoir, as one would expect. Taken together, Eqs. 1, 7, 8, 9, 10 and 11 provide a full description of the system dynamics, where the evolution of $\mu_{L}(t)$ and $T_{L}(t)$ of the finite reservoir is governed by
@@
$\dot{T}_{L}(t)$ $=\frac{\dot{Q}_{L}^{\textrm{QD}}+\dot{Q}_{L}(t)}{C_{\text{el}}\left(T_{L}(t)\right)}.$
@@ III Stationary state transport properties
$p_{1}^{s}$ $=\sum_{\alpha=L,R}\frac{\Gamma_{\alpha}^{\text{in},s}}{\Gamma_{\alpha}^{\text{in},s}+\Gamma_{\alpha}^{\text{out},s}}.$
@@
We expect our results to remain qualitatively similar for finite values of the Coulomb interaction. The stationary transport properties displayed in Fig. 2(a) would instead show a typical closed Coulomb diamond. However, the conductance lines would be modified in the same way. Similarly, the Zeeman splitting of the QD level due to a magnetic field would not alter the qualitative results.
@@
Assuming $\delta\mu_{L},\delta T_{L}$ to be small we expand Eqs. 1 and 13 to first order in the small deviations. This is a good approximation when the finite reservoir is not (significantly) depleted because $\sigma/\Gamma\gg 1$. It does not rely on a small applied bias voltage and is therefore valid beyond a strict linear response theory approach (see Appendix A for the linear-response result). The stationary state solutions for $\delta\mu_{L}^{s}$ and $\delta T_{L}^{s}$ are given in Eq. 24 in Appendix B.
@@ IV Refrigeration of the finite reservoir
Under certain conditions, the stationary transport results show a negative heat current $\dot{Q}_{L}^{\text{QD},s}$ out of the finite reservoir, see Fig. 2(b). Here we investigate under which conditions this leads to refrigeration of the finite reservoir.
The heat current flowing out of the finite reservoir as a result of electrons tunnelling into the QD is given by Eq. 12. For $V_{b}>0$, the heat current $\dot{Q}_{L}^{\text{QD},s}$ is negative when $\epsilon>\mu_{L}^{s}$, because the corresponding electron current $I_{L}^{\text{QD},s}<0$. Electrons above the Fermi level are then removed from the finite reservoir and the temperature $T_{L}^{s}$ is lowered. In turn, for $V_{b}<0$, $I_{L}^{\text{QD},s}>0$ and $\dot{Q}_{L}^{\text{QD},s}~{}<~{}0$ when $\epsilon<\mu_{L}^{s}$, electrons below the Fermi energy $\mu_{L}^{s}$ are added, again leading to a lower temperature. Achieving cooling also requires the magnitude of the (positive) heat current flowing into the finite reservoir from the infinite reservoir to be smaller than the heat current flowing out to the QD, i.e., $\lvert\dot{Q}_{L}^{s}\rvert=\lvert-\kappa_{\textrm{el}}\Delta T^{s}+\sigma\left(\Delta\mu^{s}\right)^{2}/2\rvert<\lvert\dot{Q}_{L}^{\text{QD},s}\rvert$. If those requirements are fulfilled, a temperature $T_{L}^{s}~{}<~{}T$ is induced and the finite reservoir is cooled.

### v2

@@ Abstract
We investigate non-equilibrium transport of charge and heat through an interacting quantum dot coupled to a finite electron reservoir. Both the quantum dot and the finite reservoir are coupled to conventional electric contacts, i.e., infinite electron reservoirs, between which a bias voltage can be applied. We develop a phenomenological description of the system, combining a rate equation for transport through the quantum dot with standard expressions for bulk transport between the finite and infinite reservoirs. The finite reservoir is assumed to be in a quasi-equilibrium state with time-dependent chemical potential and temperature which we solve for self-consistently. We show that the finite reservoir can have a large impact on the stationary state transport properties, including a shift and broadening of the Coulomb diamond edges. We also demonstrate that there is a region around the conductance lines where a heat current flows out of the finite reservoir. Our results reveal the dependence of the temperature that can be reached by this thermoelectric cooling on the system parameters, in particular the coupling between the finite and infinite reservoirs and additional heat currents induced by electron-phonon couplings, and can thus serve as a guide to experiments on quantum dot-enabled thermoelectric cooling of finite electron reservoirs. Finally, we study the full dynamics of the system, with a particular focus on the timescales involved in the thermoelectric cooling.
@@ I Introduction
Experimental advances in the fabrication of nanoscale electronic devices have made it possible to employ measurements of non-equilibrium transport to deduce the properties of a (quantum) system. Such devices include a wide range of systems, e.g., molecular junctions, carbon nanotubes, graphene structures and quantum dots (QDs)[1, 2, 3, 4, 5]. They are a fundamental building block for applications such as transistors, (thermal) diodes, nanoscale heat engines, and sensors [6, 7, 8, 9, 10, 11, 12, 13]. From a theoretical point of view, the description of open quantum systems out of equilibrium remains a challenging task [14]. In non-equilibrium transport scenarios, it is common to assume that the system is in contact with infinite reservoirs at different temperatures and chemical potentials. Although this is a suitable description for a number of experimental situations, accounting for the finite aspect of reservoirs is needed in many nanoscale, non-equilibrium scenarios [15]. Various experiments on, e.g., superconducting [16, 17] or ultracold atom platforms [18] demonstrate intricate control over the non-equilibrium properties of the system, including access to the state of the reservoir [19, 20].
@@
Particle transport also induces the exchange of energy and heat between different parts of a system. It is natural to consider thermal properties and thermoelectric behavior of finite reservoirs, which are the subject of ongoing interest [34]. Generally, thermoelectric effects describe a system’s or material’s ability to convert a temperature difference to an electric voltage (or vice-versa). A high thermoelectric efficiency results from strong energy asymmetries in electron transport properties [35, 36, 8], which result from sharp resonances in meso- and nanoscale systems  [37, 38, 39]. They readily occur in QD devices, promoting QDs to an ideal platform for experimental studies of thermoelectric behavior in quantum devices, see e.g. [40, 41, 42, 43, 13, 44]. Proposals to boost the thermoelectric response of quantum systems further include utilizing non-linear [45, 46] and coherence effects [47, 48], or topological states [49].
An immediate use of thermoelectric properties is Peltier cooling, where a bias voltage can be used to force heat to flow from a cold to a hot reservoir. Such electronic cooling has been realized experimentally using QD systems [50], molecular junctions [51], normal metal-insulator-superconductor junctions [52, 53] and has been investigated in cold atomic gases [54, 55]. In this context, given that any physical device or system being refrigerated is necessarily of a finite size, it is essential to understand transport and thermoelectric behaviors in non-equilibrium settings with finite reservoirs.
@@
The paper is organized as follows. In Section II, the system is introduced, and its non-equilibrium dynamics is modeled using a combination of a rate equation for the QD and standard expressions for bulk systems for transport between the finite and infinite reservoirs. In Section III, we solve for the non-equilibrium stationary state and analyse the impact of the finite reservoir on the transport through the QD. The thermoelectric refrigeration of the finite reservoir is discussed in Section IV. Finally, we investigate the system’s full dynamics in Section V.
@@ II Model
The system consists of a single level, spin-degenerate QD tunnel-coupled to an infinite reservoir on the right and a finite reservoir on the left, see Fig. 1. The finite reservoir is additionally in contact with a second infinite reservoir to the left. All reservoirs are described by non-interacting electrons. For the finite reservoir we assume that the internal relaxation is fast compared to any other time-scale in the system, a common assumption for bulk metals and semiconductors. Then, it relaxes to the quasi-equilibrium state with a well-defined (but time-dependent) chemical potential $\mu_{L}(t)$ and temperature $T_{L}(t)$, and is described by the Fermi-Dirac distribution $f(E,\mu_{L}(t),T_{L}(t)$ with $f(E,\mu,T)~{}=~{}\{\exp\left[(E-\mu)/T\right]+1\}^{-1}$. Throughout the paper we set Boltzmann’s constant $k_{B}=1$, the elementary charge $e=1$ and $\hbar=1$.
@@
where $\Gamma_{L}$ is the tunneling rate between the QD and finite reservoir and $\epsilon$ is the energy of the QD level. The total rate $\Gamma_{L}^{\text{in}}\sim 2\Gamma_{L}$ in Eq. 2 takes into account the spin degeneracy of the state with one electron on the QD. For the exchange with the right reservoir the tunnel rates are time-independent and given by
@@
$J^{\text{QD}}_{L/R}(t)$ $=\epsilon I^{\text{QD}}_{L/R}(t).$
@@
where $\tilde{T}(t)=(T_{L}(t)+\bar{T}_{L})/2$. In the limit where the difference in temperature of the left infinite and finite reservoir is negligible, Eq. 6 reduces to the Wiedemann-Franz law [58].
@@
Here, we introduced the heat currents of the finite reservoir
@@
$\dot{Q}_{L}(t)$ $=J_{L}(t)-\mu_{L}(t)I_{L}(t),$
which are positive for heat flowing into the reservoir and negative for heat flowing out of it. This shows that the dynamics of the reservoir temperature $T_{L}(t)$ is governed by the heat flows into the reservoir, as one would expect. Taken together, Eqs. 1, 7, 8, 9, 10 and 11 provide a full description of the system dynamics for arbitrary applied bias voltages and temperatures $\bar{T}_{\alpha}\gg\Gamma_{\alpha}$. In particular, the evolution of $\mu_{L}(t)$ and $T_{L}(t)$ of the finite reservoir is governed by
@@
$\dot{T}_{L}(t)$ $=\frac{\dot{Q}_{L}^{\textrm{QD}}(t)+\dot{Q}_{L}(t)}{C_{\text{el}}\left(T_{L}(t)\right)}.$
@@ III Stationary state transport properties
$p_{1}^{s}$ $=\frac{\sum_{\alpha}\Gamma_{\alpha}^{\text{in},s}}{\sum_{\alpha}\left(\Gamma_{\alpha}^{\text{in},s}+\Gamma_{\alpha}^{\text{out},s}\right)}.$
@@
Assuming $\delta\mu_{L},\delta T_{L}$ to be small we expand Eqs. 1 and 13 to first order in the small deviations. This is a good approximation when the finite reservoir is not (significantly) depleted because $\sigma/\Gamma\gg 1$. It does not rely on a small applied bias voltage (see Appendix A for the linear-response result for small bias voltage across the QD). The stationary state solutions for $\delta\mu_{L}^{s}$ and $\delta T_{L}^{s}$ are given in Eq. 24 in Appendix B.
@@ IV Refrigeration of the finite reservoir
Cooling takes place when the net heat current is negative, which is equivalent to heat flowing out of the finite reservoir. The heat current consists of two contributions defined in Eq. 12, $\dot{Q}_{L}^{\text{QD},s}$ transported due to electrons tunneling into (out of) the QD and $\dot{Q}_{L}^{s}$ flowing between the finite and infinite reservoirs. Here we investigate the conditions necessary for refrigeration of the finite reservoir.
For $V_{b}>0$ we have that $I_{L}^{\text{QD},s}<0$ and $\dot{Q}_{L}^{\text{QD},s}$ is negative when $\epsilon>\mu_{L}^{s}$. For $V_{b}<0$ we instead have that $I_{L}^{\text{QD},s}>0$ and $\dot{Q}_{L}^{\text{QD},s}$ is negative when $\epsilon<\mu_{L}^{s}$. The heat current $\dot{Q}_{L}^{s}$ from the infinite reservoir will always positive if $T_{L}^{s}<T$. Therefore, the magnitude of $\dot{Q}_{L}^{s}$ has to be smaller than the heat current flowing out to the QD, i.e., $\lvert\dot{Q}_{L}^{s}\rvert=\lvert-\kappa_{\textrm{el}}\Delta T^{s}+\sigma\left(\Delta\mu^{s}\right)^{2}/2\rvert<\lvert\dot{Q}_{L}^{\text{QD},s}\rvert$. If those requirements are fulfilled the net heat current is negative and a temperature of the finite reservoir $T_{L}^{s}~{}<~{}T$ is induced.